find the minimum force required to keep a book up. What is the maximum allowable force, as a function of θ and μ? Is there a special angle that arises? Given μ, make a rough plot of the allowed values of F for −π/2 < θ < π/2.
Answer #1

(a) Find the minimum force required to keep a book up.

Fmin = N = mg sin \theta

(b) What is the maximum allowable force, as a function of \theta and \mu?

F = fs,max

Fmax = \mus FN

Fmax = \mus mg cos \theta

(c) Is there a special angle that arises?

No